# An Agnostic Biosignature Based on Modeling Panspermia and Terraformation

### v2

@@ 2.2.2 Mantel Test
The Mantel test666While originally developed for biostatistics, and often applied to ecological data, we use the Mantel test here because it is a straightforward and established way to measure the correlation between a distance matrix representing space, and one representing an arbitrary attribute. The Mantel test has both critics and advocates (Legendre et al., 2015; Somers & Jackson, 2022) , which we mention simply to make clear that the approach we describe is not wedded to using the Mantel test—we would expect similar results when using other distance-based tests. is a measure of correlation between two distance matrices. The resulting value is called the Mantel coefficient, and is reported alongside a p-value calculated from an approximate permutation test (indicating the proportion of randomly permuted distance matrices which have correlations greater or equal to the correlation between the non-permuted distance matrices). Simply put, the p-value indicates how unlikely this correlation is compared to random permutations of the data. We wrote the Mantel test in Julia, with the algorithm and code adapted from Python’s scikit-bio (The scikit-bio development team, 2020; Rideout et al., 2023). Here we used the Mantel test to measure the Pearson correlation between a distance matrix of all planet positions, and a distance matrix of all planet compositions. For each Mantel coefficient calculation, the p-value was generated using permutations with a 2-sided alternative hypothesis. This is approximately equal to confidence. Because the p-value quantifies how anomalous an observed composition/position association is, given the assumptions of the model that correlations should only occur from panspermia and terraformation, we treat the p-value as one measure of confidence in the space containing a biosignature. To get an idea of how sensitive the Mantel coefficient corresponding to a detection is, we plotted how it varies based on number of planets observed (Fig. A3). We find that the sensitivity of the Mantel coefficient to number of planets observed decreases exponentially, and 1000 planets seems like a reasonable choice to reflect the balance of the challenge of realistically observing planets, with the need for those planets to exhibit potentially small correlations in composition-position space. The exact shape of this plot will vary by model parameters, but is especially dependent on the distribution of planet compositions and positions (e.g., planets being evenly distributed in composition or position space, vs. extremely heterogeneous).
2.2.3 Clustering
@@
We clustered planets in each iteration of the simulation based only on their compositions, using the DBSCAN algorithm implemented in R (Hahsler et al., 2019). Briefly, compared to other clustering algorithms, DBSCAN does not require the number of clusters to be predefined, can identify clusters of varying shapes and sizes based on the density of data points in the feature space, and separates points as belonging to clusters or noise. Arguments were chosen based on advice in the documentation, with (the dimensionality of the data + 1), and eps was chosen dynamically at each iteration based on the location of the elbow in the k-nearest neighbor distance plot. Because the location of the elbow is not always obvious, we used the R implementation of the Kneedle algorithm (Satopaa et al., 2011), with , empirically determined based on visual examination of the elbow placement on the nearest neighbor curves (Fig. A4). Note that the sensitivity should be adjusted to the specific data being analyzed, and in our case the data changes at each time step. This results in the Kneedle sensitivity being appropriate for only part of the time steps in our simulation. We chose to focus on the early steps of the simulation, but the sensitivity should be adjusted if finding the curves’ elbows in later time steps. DBSCAN classifies each point (planet) as either being part of a particular cluster, or noise (meaning it does not meet the criteria to fall into a cluster based on our chosen arguments).
2.2.4 Selecting clusters
@@
We created a threshold for the spatial spread of the planets in the clusters by taking the average of the interquartile range (IQR, where the middle 50% of planets fall) of all planets across each of the dimensions, , within the spatial extent of a cube with sides of length 1/2 the size of the model space (i.e., with a volume = the model space). For the simulation analyzed in this paper, . Thus, any cluster with meets our threshold for being spatially localized777Though the extent we chose for an IQR threshold is arbitrary, it reflects the relative size of our model space and the presumption that looking for life that is still relatively spatially localized is of greater interest than looking for life which has already spread over the galaxy (the latter implying there might be other easier ways to look for life). This choice, like others, could be modified depending on other assumptions or objectives. .
@@ 2.2.5 Evaluating clustering
Sensitivity is the proportion of all terraformed planets correctly selected by the cluster, . This is in our example.
Specificity is the proportion of all non-terraformed planets correctly not selected by the cluster, . This is in our example.
Accuracy is the proportion of all planets correctly classified, . This is in our example.
@@ 2.3 Software and Availability
3 Results
3.1 Panspermia can increase the correlation between planets’ compositions and positions
We find that panspermia causes uncorrelated planets to become positively correlated in position-composition distance space, as quantified by the Mantel test (see Methods. 2.2.2).
We initialized a simulation with the intention of making terraformation easy and reliable, in order to check if a positive Mantel correlation coefficient was observable in a “best case” scenario. The simulation begins with a single terraformed planet, targets the most compositionally similar planet within ( of the maximum possible distance between planets in our model), and with terraformation keeping 1 of the 10 compositional elements from the pre-terraformed planet. This means that over time, all planets in the simulation acquire similar, but not identical compositions. We found that the Mantel coefficient indeed increased as the ratio of terraformed planets increased, until reaching a peak around a terraformed ratio of , and then decreased (Fig. 2, black). Because the Mantel coefficient is measuring the correlation between positional and compositional distance matrices of planets, this correlation will decrease if either distance matrix becomes too homogeneous. This is exactly what begins to happen with the compositional distance matrix, with the decline especially pronounced if there is perfect replication (see Appendix Fig. A1, and Appendix Sec. A for a discussion on the effect of varying simulation parameters).
While the increase in Mantel correlation coefficient we can observe is striking, what we are really interested in is how anomalous the observed correlations in position/composition space are, and not the absolute value of them. That is, we expect life spreading via panspermia and terraformation to cause not just positive mantel correlation coefficients, but anomalously high ones compared to other possible permutations of the data. As seen in Fig. 2 (in red), the p-value of our Mantel test indeed decreases as planets become terraformed, and after of planets are terraformed, (which is the maximal precision we can reach with 99 permutations).
Based on our postulates, this indicates that the Mantel test—and more specifically its p-value (which remains low even after the coefficient decreases)—can identify the presence of panspermia and terraformation, provided there are enough planets in our observed population terraformed. When we investigated how sensitive the mantel coefficient is to number of planets observed (independent of the number terraformed), we find that the value of the mantel coefficient corresponding to a similarly confident detection increases exponentially with fewer planets observed (i.e., to reach the same confidence with fewer observations, we need a higher Mantel coefficient. See Fig. A3).
@@ 3.1 Panspermia can increase the correlation between planets’ compositions and positions
3.2 Likely terraformed planets can be identified from clustering
We next attempted to select planets with high potential for having undergone terraformation. Based on the idea that terraformed planets are likely to have similar compositions, we start by clustering planets based on only their compositions, at each simulation iteration (see Methods 2.2.3). The key question is—of the clusters identified, which ones likely contain terraformed planets? To address this, we selected clusters (without using the ground truth labels of which planets were terraformed) by looking at the effect they have, when removed, on the residual space’s Mantel coefficient (Mantel contribution ; Fig. 3). We further selected clusters based on their planets’ spatial localization (IQR 25.2; Fig. 4).
@@ 3.2 Likely terraformed planets can be identified from clustering
After selecting clusters, we evaluate if they actually contain terraformed planets (Figs. 7, 8). We find that across the full range of terraformed ratios, they have extremely high specificity—close to 1.0—correctly rejecting non-terraformed planets as being non-terraformed. On the other hand, detected cluster sensitivity—a measure of correctly detecting terraformed planets as being terraformed—ranges from 1.0, down to near 0.0. Perhaps surprisingly, but encouragingly, the highest sensitivity clusters appear when fewer planets are terraformed. This is could be explained by the fact that throughout the simulation clusters remain small (in terms of number of planets), and the rest of the planets often get classified as noise (this being due to the sensitivity to elbow selection). As a consequence of decreasing sensitivity, accuracy—the proportion of all planets correctly classified—also decreases.
@@ 4 Summary and Discussion
We showed how a positive Mantel correlation coefficient with a low p-value can be used to identify when a region of space contains life, under the postulates that life can spread between planetary systems and terraform planets. Going beyond (or below) a population-scale biosignature, we demonstrated how to select, from a population of planets, clusters which have high-likelihood to contain terraformed planets. More importantly, these selected clusters have a low-likelihood to contain non-terraformed planets. This is increasingly important as the field of astrobiology reconciles with unavoidable false positives for more traditional biosignatures.
@@
We assumed that the background non-terraformed distribution of planetary compositions is uncorrelated and drawn from a flat random distribution of compositions (a reasonable prior given the high sensitivity of planet formation to initial conditions of the planetary disk; (Tasker et al., 2020; Pacetti et al., 2022)). In the event that underlying abiotic planets are correlated in position and composition space (perhaps from some undiscovered astrophysical phenomena), then the value of the Mantel correlation coefficient considered to be a statistically significant biosignature would have to increase. It’s possible that if the underlying coefficient of non-terraformed planets were too high, the spread of life would actually increase the Mantel p-value, even early in the simulation. If planetary compositions are actually clustered around a few characteristic planet types, we could represent the initial non-terraformed planets as being drawn from a distribution punctuated with narrow Gaussians in composition space, and the results might be different yet. Our results underscore the importance of better understanding the baseline abiotic diversity of planetary populations, including better constraining which features of planet formation best predict planetary characteristics over time. Our model demonstrates that such information could be used to aid life detection in the absence of making progress on a theory of living systems.
Applying our approach to observed astronomical data will also be affected by relative stellar velocities. Stars move relative to each other fast enough that, for example, the Sun’s current nearest neighbor, Alpha Centauri, will be on the opposite side of the galaxy on time scales on the order of 100 My. Naturally this is a problem if we wish to detect potential correlations on the timescale of the evolution of planets. In the worst case scenario, the approach we provide would be applicable for stars in similar phase space (Lin & Loeb, 2015), or for life which spreads much faster than relative stellar velocities—potentially limiting its applicability to life spread via directed panspermia, or via lithopanspermia on exceptionally fast moving interstellar objects.
@@
Our approach here also provides an alternative to specific chemical biosignatures at the level of planets, showing the promise of detecting anomalous features at the population-scale. Critically, these anomalous features can be explained by a model with simple hypotheses about what life does, rather than what life is, and is agnostic to our ignorance of living systems as they might exist outside of Earth.
@@
This is only an introduction to the kinds of questions which could be investigated with this technique. The basic idea of taking key features of life, and understanding how they might manifest on the scale of planets leads us to imagine properties of life that may manifest across scales, like “planetary phylogenies”—that is, heredities and lineages of planets—and how they could shed light on better understanding living systems generally. Such hypothetical characteristics of planets (and life) could be investigated with our of model.
@@
Perhaps one of the biggest limitations in our model is assuming we can reliably map observable characteristics of a planet to something as comprehensible and malleable as a vector of real numbers. For example, how the chemistry of the atmosphere, biosphere, geosphere, etc. could be reflected via the atmosphere of a planet, and what the mappings are from observed atmospheric properties (the planet’s “phenotype”) to the underlying chemistries that put constraints on the compatibility of planets (the planet’s “genotype”), and the kinds of life which can emerge on or coexist with a planet. This would further strengthen the realism of our agnostic approach with minimal assumptions, and in future work we would like to explore these ideas more.
@@
While we only walk through a simulation under a single set of parameters here, our goal was to identify a best possible case scenario for applying a technique where we assume life should be able to drive correlations in planetary position-composition space based on hallmarks of life like proliferation and environmental bi-directional feedback. Our model results show promise that life could be detected at the scale of a population of planets, using information from only 1000 (perhaps fewer) planetary atmospheres, even in the absence of any information about what kinds of planetary environments are most suitable to life, or without knowing anything about the origins of life, or the peculiarities of life’s metabolic outputs. We showed how our distributed biosignature can be further refined to detect specific clusters of terraformed planets in this space, even when they only comprise a few percent of all planets. Again, this approach does not require an independent way to “detect” life, like a smoking gun biosignature atmospheric gas. Instead this approach depends on two key assumptions about what life can do, and derives observable consequences directly from them, providing a statistical approach which can be refined by astronomical surveys, and therein lies its biggest promise.
5 Acknowledgements

### v3

@@ II.2.2 Mantel Test
The Mantel test888While originally developed for biostatistics, and often applied to ecological data, we use the Mantel test here because it is a straightforward and established way to measure the correlation between a distance matrix representing space, and one representing an arbitrary attribute. The Mantel test has both critics and advocates (Legendre et al., 2015; Somers and Jackson, 2022), which we mention simply to make clear that the approach we describe is not wedded to using the Mantel test—we would expect similar results when using other distance-based tests. is a measure of correlation between two distance matrices. The resulting value is called the Mantel coefficient, and is reported alongside a p-value calculated from an approximate permutation test (indicating the proportion of randomly permuted distance matrices which have correlations greater or equal to the correlation between the non-permuted distance matrices). Simply put, the p-value indicates how unlikely this correlation is compared to random permutations of the data. We wrote the Mantel test in Julia, with the algorithm and code adapted from Python’s scikit-bio (The scikit-bio development team, 2020; Rideout et al., 2023). Here we used the Mantel test to measure the Pearson correlation between a distance matrix of all planet positions, and a distance matrix of all planet compositions. For each Mantel coefficient calculation, the p-value was generated using permutations with a 2-sided alternative hypothesis. This is approximately equal to confidence. Because the p-value quantifies how anomalous an observed composition/position association is, given the assumptions of the model that correlations should only occur from panspermia and terraforming, we treat the p-value as one measure of confidence in the space containing a biosignature. To get an idea of how sensitive the Mantel coefficient corresponding to a detection is, we plotted how it varies based on number of planets observed (Fig. A3). We find that the sensitivity of the Mantel coefficient to number of planets observed decreases exponentially, and 1000 planets seems like a reasonable choice to reflect the balance of the challenge of realistically observing planets, with the need for those planets to exhibit potentially small correlations in composition-position space. The exact shape of this plot will vary by model parameters, but is especially dependent on the distribution of planet compositions and positions (e.g., planets being evenly distributed in composition or position space, vs. extremely heterogeneous).
II.2.3 Clustering
@@
We clustered planets in each iteration of the simulation based only on their compositions, using the DBSCAN algorithm implemented in R (Hahsler et al., 2019). Briefly, compared to other clustering algorithms, DBSCAN does not require the number of clusters to be predefined, can identify clusters of varying shapes and sizes based on the density of data points in the feature space, and separates points as belonging to clusters or noise. Arguments were chosen based on advice in the documentation, with (the dimensionality of the data + 1), and eps was chosen dynamically at each iteration based on the location of the elbow in the k-nearest neighbor distance plot. Because the location of the elbow is not always obvious, we used the R implementation of the Kneedle algorithm (Satopaa et al., 2011), with , empirically determined based on visual examination of the elbow placement on the nearest neighbor curves (Fig. A4). Note that the sensitivity should be adjusted to the specific data being analyzed, and in our case the data changes at each time step. This results in the Kneedle sensitivity being appropriate for only part of the time steps in our simulation. We chose to focus on the early steps of the simulation (because the early steps of the simulation represent the time when life is not ubiquitous in the galaxy and thus expected to be harder to find), but the sensitivity should be adjusted if finding the curves’ elbows in later time steps. DBSCAN classifies each point (planet) as either being part of a particular cluster, or noise (meaning it does not meet the criteria to fall into a cluster based on our chosen arguments).
II.2.4 Selecting clusters
@@
We created a threshold for the spatial spread of the planets in the clusters by taking the average of the interquartile range (IQR, where the middle 50% of planets fall) of all planets across each of the dimensions, , within the spatial extent of a cube with sides of length 1/2 the size of the model space (i.e., with a volume = the model space). For the simulation analyzed in this paper, . Thus, any cluster with meets our threshold for being spatially localized999Though the extent we chose for an IQR threshold is arbitrary, it reflects the relative size of our model space and the presumption that looking for life that is still relatively spatially localized is of greater interest than looking for life which has already spread over the galaxy (the latter implying there might be other easier ways to look for life). This choice, like others, could be modified depending on other assumptions or objectives. .
@@ II.2.5 Evaluating clustering
Sensitivity is the proportion of all terraformed planets correctly selected by the cluster, TP/(TP+FN). This is in our example.
Specificity is the proportion of all non-terraformed planets correctly not selected by the cluster, TN/(TN+FP). This is in our example.
Accuracy is the proportion of all planets correctly classified, TP+TN/(TP+FP+TN+FN). This is in our example.
@@ II.3 Software and Availability
III Results
III.1 Panspermia can increase the correlation between planets’ compositions and positions
We find that panspermia causes uncorrelated planets to become positively correlated in position-composition distance space, as quantified by the Mantel test (see Methods. II.2.2).
We initialized a simulation with the intention of making terraforming easy and reliable, in order to check if a positive Mantel correlation coefficient is observable in a “best case” scenario. The simulation begins with a single terraformed planet, targets the most compositionally similar planet within ( of the maximum possible distance between planets in our model), and with terraforming keeping 1 of the 10 compositional elements from the pre-terraformed planet. This means that over time, all planets in the simulation acquire similar, but not identical compositions. We found that the Mantel coefficient indeed increased as the ratio of terraformed planets increased, until reaching a peak around a terraformed ratio of , and then decreased (Fig. 2, black). Because the Mantel coefficient measures the correlation between positional and compositional distance matrices of planets, this correlation will decrease if either distance matrix becomes too homogeneous. This is exactly what begins to happen with the compositional distance matrix, with the decline especially pronounced if there is perfect replication (see Appendix Fig. A1, and Appendix Sec. A for a discussion on the effect of varying simulation parameters).
While the increase in Mantel correlation coefficient we can observe is striking, what we are really interested in is how anomalous the observed correlations in position/composition space are, and not the absolute value of them. That is, we expect life spreading via panspermia and terraforming to cause not just positive mantel correlation coefficients, but anomalously high ones compared to other possible permutations of the data. As seen in Fig. 2 (in red), the p-value of our Mantel test indeed decreases as planets become terraformed, and after of planets are terraformed, (which is the maximal precision we can reach with 99 permutations).
Based on our postulates, this indicates that the Mantel test—and more specifically its p-value (which remains low even after the coefficient decreases)—can identify the presence of panspermia and terraforming, provided there are enough planets in our observed population terraformed. When we investigated how sensitive the Mantel coefficient is to number of planets observed (independent of the number terraformed), we find that the value of the Mantel coefficient corresponding to a similarly confident detection increases exponentially with fewer planets observed (i.e., to reach the same confidence with fewer observations, we need a higher Mantel coefficient. See Fig. A3).
@@ III.1 Panspermia can increase the correlation between planets’ compositions and positions
III.2 Likely terraformed planets can be identified from clustering
We next attempted to select planets with high potential for having undergone terraforming. Based on the idea that terraformed planets are likely to have similar observable compositions, we start by clustering planets based on only their compositions, at each simulation iteration (see Methods II.2.3). The key question is—of the clusters identified, which ones likely contain terraformed planets? To address this, we selected clusters (without using the ground truth labels of which planets were terraformed) by looking at the effect they have, when removed, on the residual space’s Mantel coefficient (Mantel contribution ; Fig. 3). We further selected clusters based on their planets’ spatial localization (IQR 25.2; Fig. 4).
@@ III.2 Likely terraformed planets can be identified from clustering
After selecting clusters, we evaluate if they actually contain terraformed planets (Figs. 7, 8). We find that across the full range of terraformed ratios, they have extremely high specificity—close to 1.0—correctly rejecting non-terraformed planets as being non-terraformed. On the other hand, detected cluster sensitivity—a measure of correctly detecting terraformed planets as being terraformed—ranges from 1.0, down to near 0.0. The highest sensitivity clusters appear when fewer planets are terraformed (which is encouraging for the prospects of life detection via our method if we exist at a time when the universe is not yet saturated with life). This could be explained by the fact that throughout the simulation clusters remain small (in terms of number of planets), and the rest of the planets often get classified as noise (this being due to the sensitivity to elbow selection). As a consequence of decreasing sensitivity, accuracy—the proportion of all planets correctly classified—also decreases.
@@ IV Summary and Discussion
We show how a positive Mantel correlation coefficient with a low p-value can be used to identify when a region of space contains life, under the postulates that life can spread between planetary systems and terraform planets (whether intentionally or not). Going beyond (or below) a population-scale biosignature, we demonstrate how to select, from a population of planets, clusters which have high-likelihood to contain terraformed planets. More importantly, these selected clusters have a low-likelihood to contain non-terraformed planets. This is increasingly important as the field of astrobiology reconciles with unavoidable false positives for more traditional biosignatures. In the scenario where our galaxy is already becoming saturated with terraformed life, our detection method becomes less sensitive. Our work has several limitations and space for improvements.
@@
We assumed that the background non-terraformed distribution of planetary compositions is uncorrelated and drawn from a flat random distribution of compositions (a reasonable prior given the high sensitivity of planet formation to initial conditions of the planetary disk, Tasker et al. 2020; Pacetti et al. 2022). In the event that abiotic planets are correlated in position and composition space (perhaps from some undiscovered astrophysical phenomena), then the value of the Mantel correlation coefficient considered to be a statistically significant biosignature would have to increase. It’s possible that if the underlying coefficient of non-terraformed planets were too high, the spread of life would actually increase the Mantel p-value, even early in the simulation. If planetary compositions are actually clustered around a few characteristic planet types, we could represent the initial non-terraformed planets as being drawn from a distribution punctuated with narrow Gaussians in composition space, and the results might be different yet. Our results underscore the importance of better understanding the baseline abiotic diversity of planetary populations, including better constraining which features of planet formation best predict planetary characteristics over time. Our model demonstrates that such information could be used to aid life detection in the absence of making progress on a theory of living systems.
Applying our approach to observed astronomical data will also be affected by relative stellar velocities. Naturally this is relevant if we wish to detect potential correlations on the timescale of the evolution of planets. The relative velocity dispersion of stars depends on a variety of factors (such as temperature and age), but in the solar neighborhood generally fall within km/s (Dehnen and Binney, 1998; Fatheddin and Sajadian, 2023). This implies that pairwise distance relationships among stars within a region of scale become significantly reorganized on a timescale with an upper bound of order . Using a value of corresponding to the solar neighborhood pc (Smart et al., 2021; Golovin et al., 2023) implies My. At these timescales, velocity dispersion should dominate the mixing process, when compared to the effects of velocity shear in the Milky Way (Binney and Tremaine, 2011).
Back of the envelope calculations indicate that the potential for panspermia in this time frame may not be so dire with some improvements to spacecraft speed. There are 5200 stars within 25pc of the sun, according to the latest Catalogue of Nearby Stars (CNS5)(Golovin et al., 2023). This corresponds to a density of 0.08 stars per pc3, or 1 star per 12.5 pc3, and thus an average distance between stars of 3 pc. Assuming as we do in our model that panspermia proceeded exponentially, and in parallel, it would take only generations of travel to terraform 40 stars (which is the number we identified in our results as the minimum required for detectability). Six generations corresponds to a minimum of 18 pc (3 pc per generation to the nearest neighboring stars) that have to be traveled sequentially. Benchmarking this against the fastest man made object on trajectory to escape the solar system (the Voyager spacecraft), and the fastest detected interstellar object (comet 3I/ATLAS): terraforming 40 star systems would take My years at Voyager speeds ( km/s or pc/yr) (Cook and Brown, 2024), or years at 3I/ATLAS speeds ( km/s or pc/yr) (Team, 2026) (assuming, as we do in our model, that this also includes the time it takes to terraform). These are long timescales for humans, but within the 5 My upper bound calculated above. Moreover, there are no physical laws preventing the launch of much faster spacecraft (within the speed of light), only technological and economic ones. For example, the Breakthrough Starshot program had planned to create spacecraft that could be accelerated up to speeds of 1000 km/s. If it were possible to engage in panspermia with such spacecraft, the timescale could be reduced to a mere 18,000 years.
In the absolute worst case scenario, the approach we provide would be applicable for stars in similar phase space (Lin and Loeb, 2015), and of course for life which spreads on average faster than relative stellar velocities—potentially limiting its applicability to life spread via directed panspermia, or via lithopanspermia on exceptionally abundant and fast moving interstellar objects (however, this might cause other challenges in e.g., gravitational capture and impact survival, Totani 2023). Overall, stellar and galactic dynamics paint a complicated picture and understanding their impact on our results in more detail is a subject we aim to examine in future work.
@@
Our approach also provides an alternative to specific chemical biosignatures at the level of planets, showing the promise of detecting anomalous features at the population-scale. Critically, these anomalous features can be explained by a model with simple hypotheses about what life does, rather than what life is, and is agnostic to our ignorance of living systems as they might exist outside of Earth.
By developing a life detection approach that does not require us to assume the existence of specific bio or technosignatures, we can overcome potential issues relevant to the Fermi Paradox. Simple percolation models predict that once a civilization begins the process of galactic expansion, it can proceed to rapidly spread throughout the entire galaxy. Recent work demonstrates that this is contingent on other factors like civilization lifetimes, and it’s possible to end up in a steady state where not all planets with the potential to support life will actually contain life, even after settlement fronts have spread across all of galactic space (Carroll-Nellenback et al., 2019). We might expect to have already detected life if we were currently living in such scenarios. One reason this might not be the case is if technology that life is using is too advanced for us to recognize. But another intriguing solution is that life eventually becomes so sustainable that it becomes indistinguishable from a natural non-intelligent or non-biotic planet (Schroeder, 2011; Likavčan, 2025). The method we proposed here could nevertheless be able to detect interstellar life in either of these scenarios.
@@
This discussion is only an introduction to the kinds of questions which could be investigated with this technique. The basic idea of taking key features of life, and understanding how they might manifest on the scale of planets leads us to imagine properties of life that may manifest across scales, like “planetary phylogenies”—that is, heredities and lineages of planets—and how they could shed light on better understanding living systems generally. Such hypothetical characteristics of planets (and life) could be investigated with our model.
@@
Perhaps one of the biggest limitations in our model is assuming we can reliably map observable characteristics of a planet to something as comprehensible and malleable as a vector of real numbers. For example, how the chemistry of the atmosphere, biosphere, geosphere, etc. could be reflected via the atmosphere of a planet, and what the mappings are from observed atmospheric properties (the planet’s “phenotype”) to the underlying chemistries that put constraints on the compatibility of planets (the planet’s “genotype”), and the kinds of life which can emerge on or coexist with a planet. This would further strengthen the realism of our agnostic approach with minimal assumptions, and in future work we would like to explore these ideas in detail.
@@
While we only walk through a simulation under a single set of parameters here, our goal is to identify a best possible case scenario for applying a technique where we assume life should be able to drive correlations in planetary position-composition space based on hallmarks of life like proliferation and environmental feedback. Our model results show promise that life could be detected at the scale of a population of planets, using information from only 1000 (perhaps fewer) planetary atmospheres, even in the absence of any information about what kinds of planetary environments are most suitable to life, or without knowing anything about the origins of life, or the peculiarities of life’s metabolic outputs. We showed how our distributed biosignature can be further refined to detect specific clusters of terraformed planets in this space, even when they only comprise a few percent of all planets. Again, this approach does not require an independent way to “detect” life, like a smoking gun biosignature atmospheric gas. Instead this approach depends on two key assumptions about what life can do, and derives observable consequences directly from them, providing a statistical approach which can be refined by astronomical surveys, and therein lies its biggest promise.
V Acknowledgments